How much heat is needed to change 5g of ice at 0 degrees C to water at 50 degrees C
1
Expert's answer
2020-05-26T12:52:04-0400

Solution.

"m=5g=5\sdot10^{-3}kg;"

"t_1=0^oC;"

"t_2=50^oC;"

"\lambda=332\sdot10^3J/kg;"

"\u0441=4200J\/kg^oC;"

Heat is needed to melt the ice:

"Q_1=\lambda m;"

"Q_1=332\sdot10^3J/kg\sdot5\sdot10^{-3}kg=1660J;"

Heat is needed to heat ice water:

"Q_2=cm\Delta t;"

"Q_2=4200J/(kg^oC)\sdot5\sdot10^{-3}kg\sdot50^oC=1050J;"

The whole amount of heat:

"Q=Q_2+Q_2;"

"Q=1660J+1050J=2710J;"

Answer: "Q=2710J."